# V2V-LLM: Vehicle-to-Vehicle Cooperative Autonomous Driving with Multi-Modal Large Language Models

### v3

@@ 10 Planning Results with Temporal Inputs
Table 11: V2V-LLM’s planning performance in V2V-QA’s V2V-split in comparison with baseline methods. L2: L2 distance error. CR: Collision rate. In each column, the best results are in boldface. and the second-best results are in underline.
11 Detailed Ablation Results
Tab. 12 shows the detailed ablation results when using only the scene-level features or only the object-level features as input to our V2V-LLM. Both types of input features contribute to the final performance, and object-level features are easier for LLM to digest. Training from scratch achieves worse performance, meaning that pre-training with LLaVA’s VQA tasks improves our V2V-LLM’s performance in V2V-QA.
@@ 11 Detailed Ablation Results
Table 12: Ablation study in V2V-QA’s V2V-split. Q1: Grounding at a reference location. Q2: Grounding behind a reference object at a location. Q3: Grounding behind a reference object in a direction. QGr: Average of grounding (Q1, Q2, and Q3). Q4: Notable object identification. Q5: Planning. P: Precision. R: Recall. L2: L2 distance error. CR: Collision rate. Comm: Communication cost.
12 Additional Dataset Statistics
For the grounding questions (Q1, Q2, Q3) and the notable object identification question (Q4), a QA pair can be categorized into either a positive case or a negative case. If at least one object exists that satisfies the condition specified in the questions, the corresponding QA pair is a positive case. Otherwise, it is a negative case. Tabs. 13 and 14 summarizes the numbers of QA pairs in each category, for V2V-QA’s V2V-split and V2X-split respectively. This table shows that V2V-QA has sufficient positive and negative data samples in both training and testing sets for each of these QA pairs. The planning task (Q5) is excluded from Tabs. 13 and 14, as each planning QA pair inherently includes a ground-truth future trajectory in its corresponding answer.
@@ 12 Additional Dataset Statistics
Table 13: Dataset statistics of our V2V-QA’s V2V-split on positive and negative samples.
@@
Table 14: Dataset statistics of our V2V-QA’s V2X-split on positive and negative samples.
We also visualize our V2V-split distribution of ground truth answer locations relative to the asking CAV for the grounding questions (Q1, Q2, Q3) and the notable object identification question (Q4), as shown in Figs. 7, 8, 9 and 10. In our coordinate system, $x$ is the CAV’s front direction, and $y$ is the CAV’s right direction. For the planning question (Q5), we show the distribution of the ending waypoints in the ground truth answer future trajectories, as shown in Fig. 11. We visualize the location distribution of V2V-QA’s V2X-splitin Figs. 12, 13, 14, 15 and 16. These figures indicate that our V2V-QA has diverse spatial distributions in the driving scenes. Compared to NuScenes [3], our  V2V-QA has larger ranges and standard deviations of the ground-truth ending waypoints, as shown in Tab. 15. Therefore, the planning task in our  V2V-QA is more challenging.
@@
Table 15: Ranges and standard deviations of ground-truth ending waypoints.
@@
13 Additional Qualitative Results
We show more qualitative results of our proposed V2V-LLM and other baseline methods in the testing set of V2V-QA’s grounding task in Figs. 17 to 20, notable object identification task in Figs. 21 to 22, and planning task in Figs 23 to 24. The baseline methods include no-fusion, early-fusion, and intermediate-fusion: AttFuse [52], V2X-ViT [51], and CoBEVT [50]. Results of V2X-split can be seen in Figs. 25 to 31. In general, our proposed  V2V-LLM’s outputs are closer to the ground-truth answers, in comparison to other baseline methods’ results.
Figure 17: V2V-LLM and baseline methods’ grounding results on V2V-QA’s V2V-split testing set. Magenta $\circ$: reference locations in questions. Yellow $+$: model output locations. Green $\circ$: ground-truth answers.
Figure 18: V2V-LLM and baseline methods’ grounding results on V2V-QA’s V2V-split testing set. Magenta $\circ$: reference locations in questions. Yellow $+$: model output locations. Green $\circ$: ground-truth answers.
Figure 19: V2V-LLM and baseline methods’ grounding results on V2V-QA’s V2V-split testing set. Magenta $\circ$: reference locations in questions. Yellow $+$: model output locations. Green $\circ$: ground-truth answers.
Figure 20: V2V-LLM and baseline methods’ grounding results on V2V-QA’s V2V-split testing set. Magenta $\circ$: reference locations in questions. Yellow $+$: model output locations. Green $\circ$: ground-truth answers.
Figure 21: V2V-LLM and baseline methods’ notable object identification results on V2V-QA’s V2V-split testing set. Magenta curve: planned future trajectories in questions. Green $\circ$: ground-truth notable object locations. Yellow $+$: model identification outputs.
@@ 13 Additional Qualitative Results
Figure 25: V2V-LLM and baseline methods’ grounding results on V2V-QA’s V2X-split testing set. Magenta $\circ$: reference locations in questions. Yellow $\times$: model output locations. Green $\circ$: ground-truth answers.
Figure 26: V2V-LLM and baseline methods’ grounding results on V2V-QA’s V2X-split testing set. Magenta $\circ$: reference locations in questions. Yellow $\times$: model output locations. Green $\circ$: ground-truth answers.
Figure 27: V2V-LLM and baseline methods’ grounding results on V2V-QA’s V2X-split testing set. Magenta $\circ$: reference locations in questions. Yellow $\times$: model output locations. Green $\circ$: ground-truth answers.
Figure 28: V2V-LLM and baseline methods’ notable object identification results on V2V-QA’s V2X-split testing set. Magenta curve: planned future trajectories in questions. Green $\circ$: ground-truth notable object locations. Yellow $\times$: model identification outputs.
@@
14 Limitation
Fig. 32 shows failure cases of V2V-LLM’s planning results on V2V-QA’s testing set. In a few frames, the model generates future trajectories in the lane of the opposite traffic direction. A potential solution and future work is to include HD map information as additional input to the model. Currently, this approach is not feasible because the base dataset V2V4Real [53] has not released its HD map to the public.

### v4

@@ IX Planning Results with Temporal Inputs
X Detailed Ablation Results
Table XV shows the detailed ablation results when using only the scene-level features or only the object-level features as input to our V2V-LLM. Both types of input features contribute to the final performance, and object-level features are easier for LLM to digest. Training from scratch achieves worse performance, meaning that pre-training with LLaVA’s VQA tasks improves our V2V-LLM’s performance in V2V-QA.
TABLE XV: Ablation study in V2V-QA’s V2V-split. Q1: Grounding at a reference location. Q2: Grounding behind a reference object at a location. Q3: Grounding behind a reference object in a direction. QGr: Average of grounding (Q1, Q2, and Q3). Q4: Notable object identification. Q5: Planning. P: Precision. R: Recall. L2: L2 distance error. CR: Collision rate. Comm: Communication cost.
@@ X Detailed Ablation Results
XI Additional Dataset Statistics
For the grounding questions (Q1, Q2, Q3) and the notable object identification question (Q4), a QA pair can be categorized into either a positive case or a negative case. If at least one object exists that satisfies the condition specified in the questions, the corresponding QA pair is a positive case. Otherwise, it is a negative case. Tables XVI and XVII summarize the numbers of QA pairs in each category, for V2V-QA’s V2V-split and V2X-split respectively. This table shows that V2V-QA has sufficient positive and negative data samples in both training and testing sets for each of these QA pairs. The planning task (Q5) is excluded from Tables XVI and XVII, as each planning QA pair inherently includes a ground-truth future trajectory in its corresponding answer.
TABLE XVI: Dataset statistics of our V2V-QA’s V2V-split on positive and negative samples.
@@ XI Additional Dataset Statistics
TABLE XVII: Dataset statistics of our V2V-QA’s V2X-split on positive and negative samples.
@@
We also visualize our V2V-split distribution of ground truth answer locations relative to the asking CAV for the grounding questions (Q1, Q2, Q3) and the notable object identification question (Q4), as shown in Figures 6, 7, 8, and 9. In our coordinate system, $x$ is the CAV’s front direction, and $y$ is the CAV’s right direction. For the planning question (Q5), we show the distribution of the ending waypoints in the ground truth answer future trajectories, as shown in 10. We visualize the location distribution of V2V-QA’s V2X-splitin Figures 11, 12, 13, 14, and 15. These figures indicate that our V2V-QA has diverse spatial distributions in the driving scenes. Compared to NuScenes [1], our  V2V-QA has larger ranges and standard deviations of the ground-truth ending waypoints, as shown in Table XVIII. Therefore, the planning task in our  V2V-QA is more challenging.
TABLE XVIII: Ranges and standard deviations of ground-truth ending waypoints.
@@
XII Additional Qualitative Results
We show more qualitative results of our proposed V2V-LLM and other baseline methods in the testing set of V2V-QA’s grounding task in Figures 16 to 19, notable object identification task in Figures. 20 to 21, and planning task in Figures 22 to 23. The baseline methods include no-fusion, early-fusion, and intermediate-fusion: AttFuse [36], V2X-ViT [35], and CoBEVT [33]. Results of V2X-split can be seen in Figs. 24 to 30. In general, our proposed  V2V-LLM’s outputs are closer to the ground-truth answers, in comparison to other baseline methods’ results.
Figure 16: V2V-LLM and baseline methods’ grounding results on V2V-QA’s V2V-split testing set. Magenta $\circ$: reference locations in questions. Yellow ++: model output locations. Green $\circ$: ground-truth answers.
Figure 17: V2V-LLM and baseline methods’ grounding results on V2V-QA’s V2V-split testing set. Magenta $\circ$: reference locations in questions. Yellow ++: model output locations. Green $\circ$: ground-truth answers.
Figure 18: V2V-LLM and baseline methods’ grounding results on V2V-QA’s V2V-split testing set. Magenta $\circ$: reference locations in questions. Yellow ++: model output locations. Green $\circ$: ground-truth answers.
Figure 19: V2V-LLM and baseline methods’ grounding results on V2V-QA’s V2V-split testing set. Magenta $\circ$: reference locations in questions. Yellow ++: model output locations. Green $\circ$: ground-truth answers.
Figure 20: V2V-LLM and baseline methods’ notable object identification results on V2V-QA’s V2V-split testing set. Magenta curve: planned future trajectories in questions. Green $\circ$: ground-truth notable object locations. Yellow ++: model identification outputs.
@@ XII Additional Qualitative Results
Figure 24: V2V-LLM and baseline methods’ grounding results on V2V-QA’s V2X-split testing set. Magenta $\circ$: reference locations in questions. Yellow ++: model output locations. Green $\circ$: ground-truth answers.
Figure 25: V2V-LLM and baseline methods’ grounding results on V2V-QA’s V2X-split testing set. Magenta $\circ$: reference locations in questions. Yellow ++: model output locations. Green $\circ$: ground-truth answers.
Figure 26: V2V-LLM and baseline methods’ grounding results on V2V-QA’s V2X-split testing set. Magenta $\circ$: reference locations in questions. Yellow ++: model output locations. Green $\circ$: ground-truth answers.
Figure 27: V2V-LLM and baseline methods’ notable object identification results on V2V-QA’s V2X-split testing set. Magenta curve: planned future trajectories in questions. Green $\circ$: ground-truth notable object locations. Yellow ++: model identification outputs.
@@
XIII Limitation
Figure 31 shows failure cases of V2V-LLM’s planning results on V2V-QA’s testing set. In a few frames, the model generates future trajectories in the lane of the opposite traffic direction. A potential solution and future work is to include HD map information as additional input to the model. Currently, this approach is not feasible because the base dataset V2V4Real [34] has not released its HD map to the public.